# mystyle backgroundcolor=mygray, leftmargin=0.45em, rightmargin=0.45em, innertopmargin=10pt, innerbottommargin=10pt, skipabove=10pt, skipbelow=10pt

### v2

@@ III-A Vulnerability Interpretation Enhancement Prompting
Despite the robust code understanding and analysis capabilities of LLMs, they still face challenges in complex reasoning tasks that require deep understanding of code, strong reasoning abilities, and specialized knowledge of vulnerabilities. These challenges manifest as insufficient detection accuracy and vague vulnerability analysis. To guide LLMs toward better understanding the goals of vulnerability explanation and to enhance their comprehension and analysis of vulnerable code, we have integrated instruction-based fine-tuning techniques with the contextual learning capabilities of LLMs. By leveraging prior knowledge of open-source code vulnerabilities, we design prompt templates for data annotation, fine-tuning, reasoning, and evaluation, effectively empowering the application of LLMs in vulnerability detection and explanation tasks across all crucial stages of the framework.
Figure 3 illustrates the design of our prompt templates. These templates consist of four main components: task description, specific instructions, generation examples, and sample inputs. (1) Task description: Provides a specific template for the current vulnerability detection and explanation, including information on the types of vulnerabilities being detected and the basic input-output format. This helps the LLM understand task requirements and grasp the background knowledge of the vulnerabilities. (2) Specific instructions: Include requirements for the LLM’s input-output format, such as output steps, the range of vulnerability types to focus on, and the length of the output. These requirements leverage the LLM’s instruction-following ability to ensure uniform and standardized output formats, facilitating subsequent content use and analysis. (3) Generation examples: Provide manually screened samples of vulnerable code snippets and effective explanation data pairs to help the annotation model better understand the task and generation goals. (4) Sample input: Used for effectively inputting the code to be explained, and during the annotation stage, it also includes corresponding labels and other relevant information such as CWE, CVE descriptions, or commit messages as supplementary inputs.
Figure 3: Prompting template of Automated Vulnerability Interpretation Labelling.
Through testing and evaluating various vulnerability detection and explanation tasks, we have verified that the prompt templates effectively achieved the expected goals in all stages, resulting in well-performing fine-tuned specialized vulnerability detection and explanation models.
@@ III-B Key Code Extraction Based on Chain-of-Thought (CoT)
One of the major challenges in explaining code vulnerabilities is extracting key code statements closely related to the target vulnerability type from long code snippets. Traditional static analysis methods extract code pattern information through abstract syntax trees, program dependence graphs, or control flow graphs. However, these methods are often difficult to apply in diverse vulnerability detection scenarios. To further enhance the ability of large models to analyze complex code structures, we propose a chain of thought (CoT) enhancement method based on key code extraction. This method leverages prompt engineering to use LLMs for automated key code extraction of vulnerable code. The extracted key code guides the fine-tuning model in locating suspicious vulnerabilities through a thought chain, thereby enabling targeted detection and explanation of the target vulnerabilities.
The extracted key statements are primarily based on the semantic information of the code itself and the type of vulnerability being focused on. After obtaining the key statements for each vulnerability data instance, these key statements are integrated into the prompts used during the model fine-tuning phase in the form of CoT. This approach guides the model step by step to reference the key statements to complete the vulnerability detection, localization, and explanation.
@@ IV-A1 Experimental Setup
Backbone LLMs. We use the Codellama [39] and Llama3 as our backbone LLMs for fine-tuning. CodeLlama is initialized with the weights of Llama2 [40] and fine-tuned on a specialized code dataset, thus possessing strong code understanding and generation capabilities. The Llama-family LLMs have been widely used by prior studies related to LLMs for software engineering [41, 42, 43]. In light of the natural language instruction comprehension capability of the Instruct version and the training cost, we selected CodeLlama-13B-Instruct [44] as our primary model for the experiments discussed in this chapter. For comparison purposes, we also considered the 7B version and the newly released Llama3-8B-Instruct in the discussion section.
Experimental Setting. To comprehensively evaluate the fine-tuned vulnerability LLMs, we design four sub-research questions:
@@
  • (RQ1.2) What impact does the inclusion of other types of vulnerability data in training have on the detection performance of the target type of vulnerability?
  • (RQ1.3) How effective is a unified model in detecting multiple types of vulnerabilities?
@@
  • (RQ1.4) How does the model perform in real-world scenarios with fewer data and more types of vulnerabilities to detect?
To evaluate the accuracy of vulnerability detection, we use Precision, Recall, and F1-Score as our evaluation metrics. For the binary classification scenario in RQ1.1, we directly use these three metrics. For the multi-class classification in RQ1.2, we use Weighted-F1 and Macro-F1. For the multi-label classification in RQ1.3, we use Micro-F1 and Macro-F1. To test the detection accuracy, we selected CodeT5 [45] and CodeBERT [46] as baselines, which have been commonly used by prior studies in software engineering [47, 48]. In particular, we fine-tune them for classification tasks by adding a linear classification layer.
@@
Implementation Details. We implement our approach using the Transformers [17] and PEFT [49] libraries with the PyTorch platform. All experiments are conducted on two NVIDIA A100-SXM4-80GB GPUs platforms, with the token length limit set to 2048. We use the AdamW optimizer and train the models for 3 epochs. For LoRA configuration, we set the learning rate to 0.0003, weight decay to 0.01, the LoRA rank is set to 16, the LoRA scaling factor to 16 and the dropout to 0.05.
@@
IV-A2 RQ1.1: Detection with specified vulnerability type using a dedicated model
@@
To evaluate the detection capability for a single type of vulnerability, we fine-tune each vulnerability type on SeVC to obtain a specialized model for that particular type. The sample size of each type is shown in Table II. In the training and inference prompts, we specify the type of vulnerability to help narrow the scope of detection. During training, we use explanations annotated by GPT-3.5, which include vulnerability location information and specific details, as the target output for the model to perform detection and explanation tasks. The inputs for the model during both training and inference are code snippets from the given examples. As a generative model, we conduct a binary classification task based on the semantics of the generated text. For vulnerability samples, the model generates corresponding explanatory information, while for non-vulnerability samples, it outputs a predefined fixed pattern (e.g., “There are no security issues”).
@@
Experimental Results. We first compared the impact of fine-tuning on the detection accuracy of CodeLlama. As shown in Table II, CodeLlama-13B-Instruct lacks precise vulnerability identification capabilities. Fine-tuning significantly improves the detection accuracy. Therefore, in subsequent tasks, we no longer compare with the original CodeLlama. From Table III we can observe that our generative model still achieves comparable performance to other fine-tuned classification models. On the one hand, this demonstrates that our method can effectively enhance the model’s understanding of this type of vulnerability, capturing the key patterns for determining its type. On the other hand, it also reflects that the detection task for a single type of vulnerability is relatively less challenging for LLMs. If the focus of vulnerability types in practice is relatively concentrated, superior performance can be achieved by fine-tuning with data of the target type.
IV-A3 RQ1.2: Detection with specified vulnerability type using an all-in-one model
To explore the impact of training with data from different types of vulnerabilities on the detection effectiveness for target vulnerabilities, we conduct experiments on SeVC by uniformly fine-tuning a model on four types of vulnerabilities and then performing individual detection for each type. Specifically, during training, we use all data from the four types, employing the same prompt as in (RQ1.1) to indicate the vulnerability type of each example for binary classification detection. During inference, we also provide the vulnerability type information for each example. The difference between RQ1.1 and RQ1.2 is that in RQ1.1, we fine-tune a dedicated model for each type of vulnerability, whereas in RQ1.2, we fine-tune an all-in-one model and use prompts to distinguish the type of vulnerability of interest.
TABLE IV: Performances Comparison between Dedicated Mode and All-in-one Mode(RQ1.2).
Accuracy Precision Recall F1
Type Dedicated All-in-one Dedicated All-in-one Dedicated All-in-one Dedicated All-in-one
API 93.0% 97.1% 91.6% 97.4% 94.5% 96.8% 93.0% 97.1%
Arithmetic 91.7% 96.6% 90.3% 97.6% 93.6% 95.6% 91.9% 96.6%
Pointer 92.7% 97.1% 93.7% 97.6% 91.2% 96.6% 92.4% 97.1%
Array 92.7% 96.6% 95.3% 97.0% 89.7% 96.3% 92.4% 96.6%
Average 92.5% 96.9% 92.7% 97.4% 92.2% 96.3% 92.4% 96.9%
Experimental Results. We present the results of the fine-tuned all-in-one model in Table IV. In terms of the performance comparison between the dedicated mode and the all-in-one mode, when given the types of vulnerabilities to be detected in the code, the all-in-one mode consistently achieves significant improvements across various metrics for all types of vulnerabilities. This indicates that the model can learn more patterns related to code vulnerabilities from data that explains multiple types of vulnerabilities, thereby enhancing its general vulnerability detection capabilities. This also reflects that increasing the data volume and the richness of the training data can further improve the performance of vulnerability-specialized LLMs. Moreover, this improvement is due to the ability to avoid confusion and interference that may arise from different types of vulnerabilities when the target type of the code to be detected is specified.
IV-A4 RQ1.3: Detection with an identification of the vulnerability type
In real software development environments, the vulnerability risks faced are often diverse, making it difficult to preemptively sense the potential types of vulnerabilities in the samples to be detected. This requires the large model to accurately identify and classify multiple types of vulnerabilities and provide specific explanations. To test the accuracy of a unified model in this context, we conduct a multi-class vulnerability detection fine-tuning on SeVC with all samples. Specifically, we add a new ‘[type]’ tag to the model’s response output to classify the vulnerability code, including non-vulnerable and the four types of vulnerabilities in SeVC. In the training and inference prompts, we no longer specify the type of vulnerability. Instead, we adopt a multi-type vulnerability detection scenario and provide specific descriptions of the four types of vulnerabilities in the task description.
TABLE V: Performances of Multi-Type Detection on SeVC(RQ1.3).
@@
Experimental Results. We present the metrics for each type for the multi-class vulnerability detection task in Table V. Based on the experimental results, we find that: (1) The non-vulnerability type code has a high precision and recall rate, indicating that the model still possesses strong vulnerability code identification capability under multi-type vulnerabilities and can effectively distinguish non-vulnerable code. (2) Compared to the detection results when the type of interest is provided, the overall performance has declined to varying degrees, indicating that the model finds it more challenging to differentiate between the types of vulnerabilities in a multi-type scenario. (3) Specifically, for each type of vulnerability, the Arithmetic Expression performs relatively better than other types. The other three types exhibit a significant degree of confusion and are difficult to identify. This analysis suggests that while the model can effectively identify non-vulnerable code, it struggles with differentiating between various types of vulnerabilities, especially when the data volume for certain types is insufficient. The results highlight the importance of balanced and representative training data for improving the model’s performance across all vulnerability types.
IV-A5 RQ1.4: Identification of the vulnerability type in datasets with more vulnerability types
In real development environments, security risks stem from a wide variety of project types and complex code structures, often involving a broader range of vulnerability models. To better resemble real development conditions, we conduct a multi-label classification task on the top 10 types of CWE in the real-world vulnerability dataset DiverseVul (where a small number of codes have multiple CWE labels). Specifically, we modify the task description to focus on CWE-type vulnerability detection and explanation tasks, adding a ‘[CWE]’ tag to generate a list of CWEs related to the target code. Additionally, we incorporate CWE descriptions as part of the output explanation to help the model better understand the meaning of each CWE type and utilize these descriptions for vulnerability analysis and identification.
TABLE VI: Performances of Multi-Label CWE Detection on DiverseVul (RQ1.4).
Metric CWE ID Overall (Micro)
@@
Experimental Results. The experimental results are presented in Table VI. From the results, we find that after fine-tuning, the specialized vulnerability models achieve good detection accuracy for the 10 types of CWE, demonstrating their ability to grasp the pattern information for each CWE type. Unlike the multi-class task in SeVC, the multi-label CWE task does not require the model to differentiate between categories. Each vulnerability type has a more fine-grained and explicit definition, avoiding potential category association issues in SeVC and resulting in better performance. This validates the feasibility of using LLMs for vulnerability detection and explanation in real software development environments, which is illustrated in the following explanation example.
@@
TABLE VII: Example of Generated Explanation on DiverseVul.
@@ IV-B RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
In this RQ, we explore the capability of the LLMVulExp in explaining the detected vulnerabilities. Although many evaluation metrics exist in the field of text generation by LLMs, there are no comprehensive criteria for the evaluation of vulnerability explanations. The explanatory information generated by the model is intended to help software developers identify and mitigate potential vulnerability risks, ensuring the accuracy of vulnerability analysis content. Additionally, it should be readable and concise, enabling developers to efficiently acquire the information and provide actionable suggestions for vulnerability remediation.
Considering the characteristics and requirements of the vulnerability explanation task, we propose three evaluation metrics: Accuracy, Clarity, and Actionability, as the evaluation criteria for vulnerability explanations. We use both manual review and automated evaluation by LLMs to assess the explanatory capabilities based on these criteria.
IV-B1 Experimental Setup
@@
In our study, we examine the explanations generated by the fine-tuned model for vulnerability code on the SeVC and DiverseVul test sets, as well as the ground truth annotations based on GPT-3.5. Due to the large scale of the dataset we use and the high evaluation cost, we randomly sample the data based on a 95% confidence level and a 5% confidence interval [50] following prior studies [51, 52, 53] for the fine-tuned model’s correctly classified instances. We obtain 384 and 224 samples from SeVC and DiverseVul, respectively. This sampling method ensures the statistical validity and representativeness of our evaluation.
@@
Our evaluation includes two parts: (1) Manual review: Two authors of this paper, both of whom have extensive research experience in software vulnerabilities and software development, independently conduct manual checks on the samples, scoring them according to the criteria. In cases of disagreement, discussions are held until a consensus is reached. The value of Cohen’s Kappa [54] in this process is 0.76, indicating substantial agreement. This rigorous manual review process ensures the reliability and accuracy of our evaluation. (2) Automated review using LLM: For the LLM automated review, we utilize GPT-3.5, which is instructed in the prompt with a detailed description of the vulnerability explanation evaluation task and each criterion. We ask the LLM to output the scores for three criteria, with 1 indicating satisfaction and 0 indicating non-satisfaction.
@@
This dual approach of manual and automated evaluation allows us to assess the quality of the explanations provided by the fine-tuned model and explore the potential of using LLMs for automated evaluation tasks, which can be valuable for scaling the evaluation process in the future.
Evaluation Metrics. The specific meanings and requirements of the three metrics are as follows:
@@
  • Accuracy: The explanation should correctly identify and describe the vulnerability, ensuring that the provided details are factually correct and relevant to the detected vulnerability.
@@
  • Clarity: The information should be presented in a clear, understandable manner, and structured in a way that facilitates easy comprehension by software developers.
@@
  • Actionability: The explanation should provide actionable suggestions for code modification and remediation, offering guidance on how to fix the identified issue.
IV-B2 Experimental Results
We present the results of our dual evaluation process in Table VIII, the suffix Gen. refers to the explanation generated by LLMVulExp and Ann. refers to the explanation annotated by GPT-3.5. All-Pos. refers to the results that meet the requirements of all of the three metrics. Combining the results from manual and LLM evaluations, we have made the following observations:
@@ IV-B2 Experimental Results
(1) Accuracy of Explanations: We find that LLMVulExp generally achieves a high Accuracy (e.g., over 90.0% for SeVC-Gen. for both the manual review and automated review) in the explanation. They essentially pinpoint the code’s vulnerability risks and specific code locations, reflecting that a key factor in the model’s explanatory power lies in the accuracy of vulnerability type identification.
(2) Clarity of Explanations: LLMVulExp achieves a clarity of 81.4% for SeVC and 94.1% for DiverseVul, respectively. The fine-tuned vulnerability explanation model demonstrates feasibility in reducing the barrier for developers to understand and analyze vulnerabilities in practical applications.
(3) Actionability of Explanations: The explanatory generated by LLMVulExp generally includes actionable modification suggestions (i.e., 93.4% for SeVC and 80.5% for DiverseVul), indicating that training for vulnerability explanation can effectively help models provide practical remediation advice for developers.
(4) Effectiveness of Annotation Method: Based on the results of SeVC-Ann. and DIV-Ann., we find that our proposed annotation method can effectively generate explanatory information for vulnerability data. It can stimulate the analytical capabilities of general-purpose large models for software vulnerabilities, reducing the subsequent annotation costs for vulnerability data.
(5) Potential of LLM in Automated Assessment: The automated assessment by the LLM is close to the capabilities of manual review. For SeVC and DiverseVul, the results of All-Pos. comparing manual review and automated review are 76.0% vs. 74.0% and 80.1% vs. 71.9%, respectively. This indicates that LLM has considerable potential in the evaluation of tasks related to the generation of vulnerability-related text. In the future, we can improve the accuracy of evaluations by using a dual verification process of manual and LLM automated reviews, while also reducing manual labor costs.
These observations suggest that the fine-tuned model and the LLM automated review process are valuable tools in the field of vulnerability detection and explanation. The model’s accuracy in identifying vulnerability types is crucial for generating high-quality explanations. The clarity of the explanations is beneficial for practical applications, making it easier for developers to understand and repair vulnerabilities. The effective annotation method can greatly reduce the cost of labeling vulnerability data, which is a significant advantage for scaling up vulnerability analysis. Furthermore, the LLM’s potential in automated assessment could lead to more efficient evaluation processes, balancing the need for accuracy with cost-effectiveness.
TABLE VIII: Distribution (%) of Vulnerability Explanation Review Results (RQ2).

### v3

@@ III-A Vulnerability Interpretation Enhancement Prompting
Despite the robust code understanding and analysis capabilities of LLMs, they face significant challenges in complex reasoning tasks that demand a deep understanding of code, advanced reasoning abilities, and specialized knowledge of vulnerabilities. These challenges often result in insufficient detection accuracy and vague vulnerability analyses. To address these limitations, we integrate instruction-based fine-tuning techniques with the contextual learning capabilities of LLMs. By leveraging prior knowledge from open-source code vulnerabilities, we design prompt templates for data annotation, fine-tuning, reasoning, and evaluation, effectively enabling the application of LLMs in vulnerability detection and explanation tasks across all critical stages of the framework.
Figure 3 illustrates the structure of our prompt templates. These templates are composed of four main components: task description, specific instructions, generation examples, and sample input. (1) Task description: Specifies the template for the current vulnerability detection and explanation task, including details about the types of vulnerabilities being addressed and the basic input-output format. This component provides the LLM with a clear understanding of task requirements and relevant background knowledge. (2) Specific instructions: Define the required input-output format for the LLM, such as step-by-step output guidelines, the range of vulnerabilities to focus on, and output length constraints. These instructions utilize the LLM’s instruction-following ability to ensure standardized and uniform outputs, which facilitates subsequent processing and analysis. (3) Generation examples: Contain manually curated samples of vulnerable code snippets paired with corresponding explanation data. These examples help the LLM better comprehend the task’s goals and improve the quality of generated outputs. (4) Sample input: Includes the code to be analyzed and, during the annotation phase, incorporates associated labels and supplementary information such as CWE types, CVE descriptions, or commit messages.
Figure 3: Prompt template of Explanation Annotation.
Through rigorous testing and evaluation across various vulnerability detection and explanation tasks, we validate that these prompt templates effectively achieve their intended objectives at all stages. The resulting fine-tuned models demonstrate improved performance in specialized vulnerability detection and explanation tasks.
@@ III-B Key Code Extraction Based on Chain-of-Thought (CoT)
One of the key challenges in explaining code vulnerabilities lies in accurately identifying and extracting critical code segments—referred to as key code—from lengthy code snippets. Key code typically includes statements, variable structures, and other components that are closely related to the vulnerability type under investigation. These elements often represent the root cause or propagation paths of vulnerabilities and provide essential context for understanding the issue. However, traditional static analysis techniques, such as abstract syntax trees, program dependence graphs, and control flow graphs, often fail to generalize effectively in diverse and complex vulnerability detection scenarios.
To address these challenges and improve the capability of large models to analyze complex code structures, we propose a novel Chain-of-Thought (CoT) enhancement method that incorporates automated key code extraction. This approach enables the model to focus on the most relevant parts of the code, thereby improving accuracy and interpretability. Using the prompt template illustrated in Figure LABEL:fig:prompt_key, we employ LLMs to automatically extract key statements by analyzing the semantic context of the code and the specific vulnerability type. These extracted statements form a structured CoT, guiding the fine-tuned model through a step-by-step process for targeted vulnerability detection, location, and explanation. For example, in a buffer overflow scenario, key code may include array declarations, bounds-checking statements, or pointer dereferences that contribute directly to the vulnerability. By integrating these extracted key statements into the prompts during fine-tuning, the model can systematically reference this crucial information, enhancing both the interpretability and precision of the vulnerability explanation process.
This CoT-based key code extraction methodology ensures that the model is not only detecting vulnerabilities more accurately but also providing detailed, context-rich explanations that developers can use to understand and address these issues effectively.
Figure 4: Prompt template of Key Code Extraction.
@@ IV-A RQ1: How effective are LLMs in detecting software vulnerabilities?
Backbone LLMs. In light of the natural language instruction comprehension capability of the Instruct version and the training cost, we selected CodeLlama-13B-Instruct [39] as our primary model for the experiments discussed in this chapter. CodeLlama is initialized with the weights of Llama2 [8] and fine-tuned on a specialized code dataset, thus possessing strong code understanding and generation capabilities. Experimental Setting. To comprehensively evaluate the fine-tuned vulnerability LLMs, we design three sub-research questions:
@@
  • (RQ1.2) How effective is the model in detecting multiple types of vulnerabilities simultaneously?
@@
  • (RQ1.3) How does the model perform in real-world scenarios with fewer data and more types of vulnerabilities to detect?
To evaluate the accuracy of vulnerability detection, we use Precision, Recall, and F1-Score as our evaluation metrics. For the binary classification scenario in RQ1.1, we directly use these three metrics. For the multi-class classification in RQ1.2, we use Weighted-F1 and Macro-F1. For the multi-label classification in RQ1.3, we use Micro-F1 and Macro-F1. To test detection accuracy, we selected CodeT5 [40] and CodeBERT [41] as baselines and fine-tuned them for classification tasks by adding a linear classification layer. These models are chosen for two primary reasons: (1) As pre-trained code models, CodeT5 and CodeBERT have demonstrated consistent and superior performance on prior vulnerability detection benchmarks [16, 14], making them reliable baselines. (2) Their inclusion allows us to assess whether fine-tuned LLMs offer an advantage over existing pre-trained code models in terms of detection accuracy. It is important to note that the focus of this study is not on developing the most optimal detection model but rather on evaluating the explanation capabilities of fine-tuned LLMs. Therefore, using these two baselines provides a meaningful comparison without detracting from the primary objective of assessing the interpretability of LLM-generated explanations.
@@
Implementation Details. We implement our approach using the Transformers[13] and PEFT[42] libraries on the PyTorch platform. All experiments are conducted on two NVIDIA A100-SXM4-80GB GPUs, with a token length limit set to 2048. The models are trained using the AdamW optimizer for 3 epochs with a batch size of 2 and data precision set to float16 to optimize memory usage and computational efficiency. We set the learning rate to 0.0003 and weight decay to 0.01. For the LoRA configuration, we set target_modules to {q_proj, k_proj, v_proj, o_proj} for SeVC and {q_proj, k_proj, v_proj, o_proj, up_proj, down_proj, gate_proj} for DiverseVul. The LoRA rank is set to 16, the LoRA scaling factor to 16, and the dropout rate to 0.05. For fine-tuning CodeBERT and CodeT5, we use the AdamW optimizer with the learning rate of 1e-5 and the weight decay of 1e-2, and train the models for 3 epochs. All other parameters are set to their default values.
@@
IV-A2 RQ1.1: Detection with specified vulnerability type using a dedicated model.
@@
To evaluate the detection capability for a single type of vulnerability, we fine-tune a separate model for each vulnerability type on SeVC, resulting in specialized models tailored to individual vulnerability types. The sample sizes of each type is shown in Table II. In both training and inference prompts, explicitly specify the target vulnerability type to narrow the scope of detection and improve precision. During training, the model utilizes explanations annotated by GPT-3.5, which include detailed information about the vulnerability location and its specific characteristics, as the target output for performing both detection and explanation tasks. The model’s input during both training and inference consists of code snippets from the corresponding dataset examples. As a generative model, the task is formulated as a binary classification based on the semantics of the generated text. For vulnerability samples, the model generates corresponding explanatory information, while for non-vulnerability samples, it produces a predefined fixed pattern (e.g., “There are no security issues”).
@@
Experimental Results. We first evaluate the impact of fine-tuning on the detection accuracy of CodeLlama. As shown in Table II, the original CodeLlama-13B-Instruct model lacks precise vulnerability identification capabilities, demonstrating suboptimal performance in detecting specific types of vulnerabilities. However, fine-tuning significantly improves detection accuracy. Consequently, we exclude the original CodeLlama from subsequent comparisons.
@@
From Table III, we observe that our generative model achieves performance comparable to other fine-tuned classification models. This result highlights two key findings: (1) Our approach effectively enhances the model’s understanding of the specific vulnerability type, enabling it to capture critical patterns for accurate detection. (2) The detection task for a single type of vulnerability is relatively less challenging for LLMs. In practical scenarios where the focus is on a limited set of vulnerability types, fine-tuning with target-specific data can yield superior performance.
IV-A3 RQ1.2: Detection with identification of multiple vulnerability types.
In real software development environments, vulnerability risks are often diverse, posing challenges in accurately predicting the specific types of vulnerabilities present in the code. This necessitates a model capable of simultaneously detecting, classifying, and explaining various vulnerability types. To evaluate the effectiveness of such a model, we train a single model on SeVC for multi-class vulnerability detection using all available samples. Specifically, a ‘[type]’ tag is appended to the model’s output to classify the detected vulnerabilities, encompassing non-vulnerable code and the four distinct types of vulnerabilities present in SeVC. Unlike the previous section, where each model is trained to detect a single type of vulnerability, here we use a unified approach to enable one model to handle all types concurrently. During training and inference, the specific vulnerability type of detected code will not appear in the prompt. Instead, a task description encompassing all four vulnerability types in the current multi-class scenario is provided, enabling the model to develop the capability to distinguish between multiple types of vulnerabilities.
TABLE IV: Performances of Multi-Type Detection on SeVC (RQ1.2).
@@
Experimental Results. The metrics for each type in the multi-class vulnerability detection task are presented in Table IV. Based on the experimental results, we observe the following: (1)The non-vulnerable code category exhibits high precision and recall, indicating that the model maintains strong capability in identifying non-vulnerable code even in a multi-type scenario. (2) Compared to detection results where the type of interest is provided, the overall performance shows varying degrees of decline, indicating that the model faces greater difficulty in distinguishing between vulnerability types. (3) Among the vulnerability types, Arithmetic Expression achieves relatively better performance, whereas the remaining three types demonstrate significant confusion, making them harder to identify. These findings indicate that, while the model is effective at recognizing non-vulnerable code, its ability to differentiate between multiple vulnerability types is limited, particularly when certain types are underrepresented in the training data. This highlights the critical role of balanced and representative training datasets in improving the model’s performance across all vulnerability types.
IV-A4 RQ1.3: Identification of a greater variety of vulnerability types.
In real-world software development environments, security risks arise from a diverse range of project types and complex code structures, often involving a broader spectrum of vulnerability categories. To better reflect these real-world conditions, we conduct a multi-label classification task using the top 10 CWE categories in the real-world vulnerability dataset DiverseVul, where a subset of the code samples is annotated with multiple CWE labels. Specifically, the task description is modified to emphasize CWE-type vulnerability detection and explanation, with the addition of a ‘[CWE]’ tag to generate a list of CWE categories relevant to the target code. Furthermore, CWE descriptions are incorporated as part of the output explanation, aiding the model in comprehending the context and semantics of each CWE type and leveraging these descriptions for vulnerability analysis and identification.
TABLE V: Performances of Multi-Label CWE Detection on DiverseVul (RQ1.3).
Metric CWE ID Overall (Micro-F1)
@@
Experimental Results. The experimental results are presented in Table V. The findings demonstrate that after fine-tuning, the specialized vulnerability models achieved high detection accuracy across the 10 CWE categories, highlighting their capability to capture the distinctive patterns associated with each CWE type. Unlike the multi-class task in SeVC, the multi-label classification in DiverseVul does not require the model to distinguish between mutually exclusive categories. Instead, each CWE type has a more granular and explicit definition, mitigating the category association challenges observed in SeVC and thereby yielding improved performance. These results validate the practicality of employing LLMs for vulnerability detection and explanation in real-world software development scenarios. The following example of a generated explanation further illustrates the model’s capacity to analyze and identify CWE-based vulnerabilities effectively.
@@
TABLE VI: Example of Generated Explanation on DiverseVul.
@@ IV-B RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
In this RQ, we explore the proficiency of the LLMVulExp in explaining the detected vulnerabilities. While numerous evaluation metrics exist for assessing text generation by LLMs [43, 44], there is currently no comprehensive framework tailored to evaluating vulnerability explanations. The explanatory content generated by the model should assist software developers in identifying and mitigating potential vulnerabilities by ensuring the accuracy of the analysis, maintaining readability and conciseness, and providing actionable suggestions for remediation.
The evaluation of explanation quality remains an open problem, as there are no definitive standards for what constitutes a ”correct” or ”complete” explanation. This inherent subjectivity poses significant challenges for effective evaluation. Therefore, the evaluation of explanations must be contextualized within the specific goals of the vulnerability explanation task. By summarizing and generalizing the “helpfulness” of the generated text, we assess explanation quality from three critical perspectives: Accuracy, Clarity, and Actionability.
IV-B1 Experimental Setup.
@@
In this study, we evaluate the explanations generated by the fine-tuned LLMVulExp for vulnerability code from the SeVC and DiverseVul test sets, alongside ground truth annotations generated using GPT-3.5. Given the large scale of the dataset and the high cost of evaluation, we employ a random sampling approach based on a 95% confidence level and a 5% confidence interval [45]. This yields 384 and 224 samples from SeVC and DiverseVul, respectively, ensuring statistical validity and representativeness in our evaluation.
@@
Our evaluation includes two parts: (1) Manual review: Two authors, with extensive research experience in software vulnerabilities and development, independently review the sampled data, scoring them based on predefined criteria. Disagreements are resolved through discussion until a consensus is reached. The Cohen’s Kappa value [46] for inter-reviewer agreement is 0.76, indicating substantial consistency and reliability. (2) Automated review using LLM: GPT-3.5 is used for automated evaluation. A detailed prompt is provided, describing the vulnerability explanation evaluation task and the specific criteria. The LLM is instructed to output scores for the three evaluation metrics, with 1 indicating satisfaction and 0 indicating non-satisfaction.
@@
This dual evaluation approach provides a robust assessment of the quality of the explanations generated by the fine-tuned model and demonstrates the feasibility of using LLMs for scalable evaluation tasks in the future.
Evaluation Metrics. The three evaluation metrics used in this study are defined as follows:
@@
  • Accuracy: The explanation should correctly identify and describe the vulnerability, ensuring the details are factually correct and relevant to the detected vulnerability.
@@
  • Clarity: The explanation should be clear, concise, and structured in a way that facilitates easy comprehension by software developers.
@@
  • Actionability: The explanation include actionable suggestions for remediation, offering practical guidance for addressing the identified vulnerability.
IV-B2 Experimental Results.
We present the results of our dual evaluation process in Table VII. The suffix Gen. refers to explanations generated by LLMVulExp, while Ann. refers to annotations from GPT-3.5. All-Pos. represents cases meeting all three evaluation metrics. Key findings include:
@@
(1) Accuracy of Explanations: LLMVulExp achieves high accuracy (e.g., over 90.0% for SeVC-Gen. in both manual and automated reviews), pinpointing vulnerability risks and code locations. This underscores the importance of accurate vulnerability type identification.
(2) Clarity of Explanations: The clarity scores are 81.4% for SeVC and 94.1% for DiverseVul, indicating the model effectively reduces cognitive barriers for developers analyzing vulnerabilities.
(3) Actionability of Explanations: Explanations from LLMVulExp include actionable suggestions for code modifications, achieving 93.4% for SeVC and 80.5% for DiverseVul. This highlights the model’s potential to support practical remediation efforts.
(4) Effectiveness of Annotation Method: Results from SeVC-Ann. and DIV-Ann. validate the effectiveness of our annotation method in generating high-quality explanatory data, reducing subsequent annotation costs.
(5) Potential of LLM in Automated Assessment: LLM-based automated evaluations closely align with manual reviews. For All-Pos., manual vs. automated results are 76.0% vs. 74.0% for SeVC and 80.1% vs. 71.9% for DiverseVul. This demonstrates LLMs’ potential to scale evaluation tasks efficiently.
While the proposed approach demonstrates strong performance, it has inherent limitations. Fully correct explanations cannot be guaranteed due to the complexity of vulnerability analysis, and manual evaluations are subject to reviewer bias. Despite these challenges, the dual-validation strategy combining manual and automated reviews offers a practical and effective solution for evaluating vulnerability explanations, balancing accuracy with scalability and cost-effectiveness.
TABLE VII: Proportion (%) of Vulnerability Explanation Review Results (RQ2).